• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 16:06:23 on October 8,2002

    Entry number 87633

    keyword=Shift Summary

    8:00 Start shift taking production data as well as parasitic d(e,e') at p_R=976.331 MeV. Fast feedback troubles - mostly off.
    8:30 Doug learns from MCC that there were beam spot problems in the injector last night. After those were corrected this morning, our rate drops by about a factor of 3.
    Last night's data are probably not of the highest quality.
    9:00 Start of ~10 minutes of target tests with higher beam current to check temperature response (Chris Keith). Eugene setting up Moller. Run 1232 still
    in progress.
    9:10 End run 1232.
    9:15 Changing HRSR momentum to 901.23 MeV.
    10:30 Moller discovers a DAQ problem which can only be fixed downstairs.
    Asking for controlled access.
    12:30 Moller DAQ repaired, although exact problem not understood. Apparently a flakey trigger module. Continuing Moller measurements.
    14:15 Noticed that left dipole and right Q1 have quenched. Dipole comes back, but not Q1R, which refuses to reset.
    14:55 Moller done. Start some brief charge asymmetry tests.
    15:00 Decide that we need to reset Q1R before doing the asymmetry test. So, we ask for access.
    15:30 Access over. Setting up for asymmetry tests. Discussion ensuing re impact on Hall B.
    16:00 Still not running at end of shift.